Please help me to solve this
problem #1
Two cars leave a town and travel in opposite directions. One car travel at the rate of 45 pkh and the other at 55 kph. In how many hours will the two cars be 350 kilometers apart?
Problem #2
How long will it takes a bus traveling at 60 kph to overtake a car traveling at 40 kph if the car had a 1.5-hour head start?
Thanks and hopping for your help
Regards,
JM

Answer by mananth(16949)

Please help me to solve this
problem #1
Two cars leave a town and travel in opposite directions. One car travel at the rate of 45 pkh and the other at 55 kph. In how many hours will the two cars be 350 kilometers apart?
one car speed = 45 kph
other car speed = 55kph
effective speed will be 55+45 = 100 kph since they are moving away from each other.
Distance = 350 km
time = distance / speed
350/100
= 3.5 hours
Problem #2
How long will it takes a bus traveling at 60 kph to overtake a car traveling at 40 kph if the car had a 1.5-hour head start?
Bus speed = 60 kph
car speed = 40 kph
effective speed = 20kph . Bus will catch up 20 kph
car has left 1.5 hours early. so in 1.5 hours it travels 60 km
Time = distance/speed = 60/20 = 3 hours bus will catch up.
m.ananth@hotmail.ca